OpenStudy (rayman5766):

Simplify the following expression. 1/3 / 3/25

OpenStudy (rayman5766):



OpenStudy (rayman5766):

what are u going to leave to

OpenStudy (helpblahblahblah):

\[\frac{ 1 }{ 3 }\div \frac{ 3 }{25}= \frac{ 1 }{ 3}\times \frac{ 25 }{ 3 }=\frac{ 8 }{ 1}\] What i did was KCF and which flips the second fractions and i estimated becasue the closest 3 can get to 25 is by dividing it by 8 and i cross out the 1 and 3... thats how i think you would do the problem